OpenStudy (anonymous):

Drey makes $10 an hour plus $15 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. Part C: Drey earned $490 in 1 week. How many hours (regular plus overtime) did he work?

OpenStudy (anonymous):

Okay so we can write $1o an hour as:\[10h\] Follow so far?

OpenStudy (anonymous):

im following thats all i had so far honestly

OpenStudy (anonymous):

We can construct an equation: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. \[W = 10x + 15o\] Where is \(x\) is hours worked, and \(o\) is overtime. Since not overtime exist we can drop that part off.\[W = 10x\] is the equation for Part A. Understand how we did that?

OpenStudy (anonymous):

yes i follow wages is the variable and 10 is per hour 15 is over time but we dont need it yet so w is now wages = 10x

OpenStudy (anonymous):

* without overtime. Remember we are not accounting for overtime since the question says no overtime is worked.

OpenStudy (anonymous):

yes i had understood that

OpenStudy (anonymous):

Okay try to attempt: Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. And show me your attempt.

OpenStudy (anonymous):

W= S$10 per hour and Y $15 per over time hour 10s+15y

OpenStudy (anonymous):

y = 40 remember you work that many hours overtime.

OpenStudy (anonymous):

okay got that now

OpenStudy (anonymous):

Great, good luck!

OpenStudy (anonymous):

wait what about part c ! lol

OpenStudy (anonymous):

Okay, well attempt it, and I will check if you are right or wrong.

OpenStudy (anonymous):

i got 70 hours 490/7

OpenStudy (anonymous):

Overtime is any hours over hour 40. So \(40 \times 10 = 400\). We still need to calculate the overtime which is $15 per hour. \(15 \times 6 = 90\) So \(40 + 6 = 46 hours\).